### v1

# Modeling Domain and Feedback Transitions for Cross-Domain Sequential Recommendation

Abstract
Nowadays, many recommender systems encompass various domains to cater to users’ diverse needs, leading to user behaviors transitioning across different domains. In fact, user behaviors across different domains reveal changes in preference toward recommended items. For instance, a shift from negative feedback to positive feedback indicates improved user satisfaction. However, existing cross-domain sequential recommendation methods typically model user interests by focusing solely on information about domain transitions, often overlooking the valuable insights provided by users’ feedback transitions. In this paper, we propose $\text{Transition}^{2}$, a novel method to model transitions across both domains and types of user feedback. Specifically, $\text{Transition}^{2}$ introduces a transition-aware graph encoder based on user history, assigning different weights to edges according to the feedback type. This enables the graph encoder to extract historical embeddings that capture the transition information between different domains and feedback types. Subsequently, we encode the user history using a cross-transition multi-head self-attention, incorporating various masks to distinguish different types of transitions. Finally, we integrate these modules to make predictions across different domains. Experimental results on two public datasets demonstrate the effectiveness of $\text{Transition}^{2}$.
Introduction
(a) An illustration of a user’s transitions: the first transition is a Type 1 transition, where the user enjoys a movie adaptation of a book, but negative (neg.) feedback arises from recommending the original book rather than its sequel; the second transition belongs to Type 2, where the user is dissatisfied with art books and recommending other types of movies results in positive (pos.) feedback.
(b) Statistical analyses on Book-Movie domains and Book-Music domains of Douban, focusing on the percentage of Type 1 transitions, Type 2 transitions, and other transitions among all cross-domain user behaviors.
Figure 1: A toy illustration and statistics of domain and feedback transitions in cross-domain sequential recommendation. Type 1 transition: when transitioning across domains, the user feedback changed from positive to negative; Type 2 transition: when the user browses the new domain, her feedback improved from negative to positive.
Traditional sequential recommendation models (Zhang et al. 2024; Kang and McAuley 2018) a user’s history within a single domain, such as music (Zhang et al. 2022) or videos (Zhao et al. 2024). However, modern recommender systems often encompass multiple domains, such as books (Anwar, Siddiqui, and Saquib Sohail 2019) and movies (Goyani and Chaurasiya 2020), leading users to transition between different domains. In this transition process, not only is the information about the domain transitions valuable (referred to domain transition), but the changes in user preference toward recommended items during domain transitions are also crucial (termed feedback transition).
We classify users’ feedback transitions during their domain transitions into two types: (1) Type 1 transition: when transitioning across domains, the user feedback changed from positive to negative; (2) Type 2 transition: when browsing the new domain, the user’s feedback improves from negative to positive. For example, as shown in Figure 1(a), for the transition of “Type 1”, a user may search for related books after watching a movie she likes. However, since the movie is adapted from a book, directly recommending the original book might be disliked because she has already watched the movie. In this case, the ideal recommendation would be the sequel to that book. For a “Type 2” transition, the user is dissatisfied with several consecutively recommended books and turns to watch movies instead. This also suggests that when a user is consistently dissatisfied with items in one domain, it may be beneficial to recommend items from another domain.
Furthermore, we analyzed the percentage of these two types of transitions among all transitions in the Douban [^1] https://www.douban.com/ data, as shown in Figure 1(b). Specifically, we analyzed the cross-domain behaviors between Book-Movie and Book-Music, focusing on the counts of “Type 1” and “Type 2” transitions. We found that these two transition types account for a significant percentage of all cross-domain behaviors ($18.31\%$ and $18.28\%$ in Book-Movie, and $14.53\%$ and $14.29\%$ in Book-Music). This finding underscores the importance of modeling cross-domain behaviors and feedback transitions. This example shows that considering negative feedback can help us better understand users’ cross-domain behavior transitions, enabling better user modeling.
Existing work typically focuses on modeling users’ domain transition behaviors, to improve the performance of cross-domain sequential recommendations. A pioneering work transfers single-domain representations learned from a single domain to other domains using a gated transfer module (Ma et al. 2019). Recent work models single-domain and cross-domain sequences separately and generates representations through a self-attention mechanism (Cao et al. 2022; Ma et al. 2024). Despite the effectiveness of existing works, they overlooked the impact of users’ feedback transitions during cross-domain behaviors, and often only focused on the positive feedback from users.
To model both the domain transitions and feedback transitions in users’ cross-domain behaviors, this paper proposes an approach named $\text{Transition}^{2}$, which models domain and feedback transitions for the cross-domain sequential recommendation. Firstly, we encode users’ mixed histories across different domains using the transition-aware graph encoder. Specifically, when constructing the graph, we connect consecutive items in the history and assign different weights to the edges based on the feedback between the two items. This allows us to capture transitions between different domains and different types of feedback. After encoding with the graph encoder, we obtain item embeddings that capture transitions between different domains and feedback types. We then input these history embeddings into a cross-transition multi-head self-attention to further model the transition information in the user’s history. To model the transitions between different domains and feedback types, we introduce different masks into the heads of the transformer’s multi-head self-attention. This allows each head to capture different types of transitions. We then fuse the outputs from different heads to obtain the final history representations which include various transition information. Finally, we use this representation for prediction tasks in different domain recommendations.
The major contributions of the paper are as follows:
  • We identified the importance of domain transitions and feedback transitions in cross-domain sequential recommendation and validated this through data statistics.
  • We proposed $\text{Transition}^{2}$, which models domain and feedback transitions for cross-domain sequential recommendations using a GNN and a transformer equipped with different masks.
  • Experiment results on two public datasets demonstrate the effectiveness of $\text{Transition}^{2}$. $\text{Transition}^{2}$ outperforms existing sequential recommendation and cross-domain sequential recommendation models.
Related Work
Cross-Domain Sequential Recommendation (CDSR) addresses the challenge of recommending items across diverse domains by leveraging various methodological approaches. Perera and Zimmerman pioneered using timestamp information to divide user interactions into temporal itemsets, enhancing the modeling of users’ dynamic interests across domains (Perera and Zimmermann 2020a, b). Zhang et al. further advanced this by proposing the CGN model, utilizing dual generators to map itemsets across domains concurrently (Zhang et al. 2020). Another strand of research focuses on domain-specific sequential modeling, exemplified by $\pi$-Net (Ma et al. 2019) and DASL (Li et al. 2021), which employ GRUs and attention mechanisms for knowledge transfer. Recent innovations like DA-GCN (Chen et al. 2019) and MIFN (Ma et al. 2022) introduce graph-based approaches to link domain-specific item sequences, while industry applications such as SEMI (Lei et al. 2021) and RecGURU (Li et al. 2022) apply adversarial learning and multi-modal data fusion for cross-domain short-video recommendations. Novel hybrid models like C2DSR (Cao et al. 2022) integrate graphical and attentional mechanisms, employing contrastive objectives to enhance both intra-domain and cross-domain user representations. DREAM (Ye, Li, and Yao 2023) focuses on modeling decoupled representations for both single- and cross-domain. In summary, current CDSR methods fail to effectively extract and integrate cross-domain transition information, disregarding user negative feedback and thus neglecting the transition of information between positive and negative feedback. These challenges are crucial in cross-domain recommendation scenarios for predicting future user interactions.
Figure 2: The overall framework of $\text{Transition}^{2}$ consists of four modules: (1) Embedding Layer for embedding initialization, (2) Transition-Aware Graph Encoder for graph representation of user sequence considering transition, (3) Cross-Transition Multi-Head Self-Attention for capturing transition information with various masks, (4) Prediction Module for optimization.
Task Formulation
In this paper, we discuss a comprehensive CDSR scenario where each user history involves two domains, namely domain $A$ and $B$. Each instance $S_{A},S_{B},S_{C}$ corresponds to a specific user. For any given instance, $S_{A}=[(a_{1},y_{1}),\ldots,(a_{|S_{A}|},y_{|S_{A}|})]$ and $S_{B}=[(b_{1},y^{\prime}_{1}),\ldots,(b_{|S_{B}|},y^{\prime}_{|S_{B}|})]$ represent the single-domain user histories, where $y_{i}$ and $y^{\prime}_{i}$ indicate the feedback associated with $a_{i}$ and $b_{i}$, respectively, taking values of $+1$ (positive feedback) or $-1$ (negative feedback). The sequence $S_{C}=[(b_{1},y^{\prime}_{1}),(a_{1},y_{1}),\ldots,(b_{|S_{B}|},y^{\prime}_{|S_{B}|}),\ldots,(a_{|S_{A}|},y_{|S_{A}|})]$ represents the cross-domain user history, formed by merging $S_{A}$ and $S_{B}$ in chronological order, where each $a\in A$ and $b\in B$ are the interacted items, and ||$\cdot$|| ⋅ | denotes the total number of items. Note that $A$ and $B$ denote the entire item sets in domain $A$ and domain $B$, respectively. Given the observed user history $(S_{A},S_{B},S_{C})_{u}$, the goal of CDSR is to predict the next item:
$\arg\max_{a_{i}\in A}P_{A}(a_{|S_{A}|+1}|S_{A},S_{B},S_{C})\quad\text{if the next item is in }A,$
$\arg\max_{b_{j}\in B}P_{B}(b_{|S_{B}|+1}|S_{A},S_{B},S_{C})\quad\text{if the next item is in }B,$
where $P_{A}(a_{i}|S_{A},S_{B},S_{C})\in\mathbb{R}^{|A|}$ and $P_{B}(b_{j}|S_{A},S_{B},S_{C})\in\mathbb{R}^{|B|}$ denote the probability of the candidate item in domain $A$ and $B$, respectively, with the highest probability item being chosen as the next recommended item.
Transition2: The Proposed Algorithm
In this section, we will introduce our proposed model $\text{Transition}^{2}$.
Overview of Transition2
The overall framework of our model Transition2 is illustrated in Figure 2. To effectively capture domain-transition and feedback-transition information: (1) $\text{Transition}^{2}$ first initializes the embeddings for two single-domain sequences and the cross-domain sequence with three embedding layers. (2) $\text{Transition}^{2}$ then constructs transition-aware graph encoder based on users’ cross-domain and single-domain sequences. Unlike previous methods that only utilize positive feedback for graph construction, we incorporate negative feedback and adjust the weight matrix of the graph neural network accordingly. (3) After obtaining the representation of the user’s historical sequences through the graph encoder, we design a cross-transition multi-head self-attention mechanism. This mechanism calculates attention between and within different behaviors of the user’s sequence to capture transition information. (4) The predicted scores are then calculated based on the final sequence representations, optimizing both single-domain and cross-domain losses.
Transition-Aware Graph Encoder
Inspired by the advantages of GNNs in handling sequential recommendations (Wu et al. 2019; Xu et al. 2019), we leverage GNNs to transform users’ sequential behaviors into graph representations. Additionally, to effectively capture domain transitions and feedback transitions, we have introduced transition-aware graph encoder.
Embedding Initialization
For the sequential recommendation settings, we construct single-domain graphs $G_{A}$ and $G_{B}$, as well as a cross-domain graph $G_{C}$, based on all users’ historical sequences, where only item nodes are included. Consequently, we initialize the item embeddings (Qu et al. 2023, 2024) in the three graphs as $\mathbf{E}^{(0)}_{A}=[\mathbf{e}^{(0)}_{A1};\mathbf{e}^{(0)}_{A2};\cdots]\in\mathbb{R}^{{|A|}\times d}$, $\mathbf{E}^{(0)}_{B}=[\mathbf{e}^{(0)}_{B1};\mathbf{e}^{(0)}_{B2};\cdots]\in\mathbb{R}^{{|B|}\times d}$, $\mathbf{E}^{(0)}_{C}=[\mathbf{e}^{(0)}_{C1};\mathbf{e}^{(0)}_{C2};\cdots]\in\mathbb{R}^{{(|A|+|B|)}\times d}$, respectively.
Transition-Aware Graph Construction
The sequence graphs are constructed based on the user’s interaction sequences by treating each item in the sequence as a node and adding edges between consecutive items. The complete graphs is obtained by combining all sequences in the dataset. Unlike previous methods that only consider items with positive feedback for graph construction, we also include items with negative feedback. The comprehensive cross-domain graph includes all items with both positive and negative feedback from domains A and B. However, this approach introduces some challenges. Firstly, during the propagation phase of the graph, nodes receive information from their neighbors. Essentially, items with different feedback should propagate opposite information. For example, items with positive feedback should receive positive information from neighboring items with positive feedback and negative information from neighboring items with negative feedback. Therefore, we incorporate feedback-transition information to obtain the transition matrix $\mathbf{T}$. If two adjacent nodes have different types of feedback, we set their corresponding values opposite. Specifically, for two adjacent items $i$ and $j$ with corresponding feedback $y_{i}$ and $y_{j}$, the transition matrix is defined as follows:
$\textbf{T}_{ij}=\begin{cases}1&\text{if }y_{i}=y_{j},\\ -1&\text{if }y_{i}\neq y_{j},\\ 0&\text{if }i\text{ and }j\text{ are not adjacent}.\end{cases}$ (1)
We then obtain the transition-aware adjacency matrix of the item-item transition graph as
$\mathbf{W}=\left(\begin{array}[]{cc}\mathbf{0}&\mathbf{T}\\ \mathbf{T}^{\top}&\mathbf{0}\\ \end{array}\right).$ (2)
To stabilize training, we use the normalized form:
$\mathbf{\widehat{W}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{W}\mathbf{D}^{-\frac{1}{2}}.$ (3)
In this equation, $\mathbf{D}\in\mathbb{R}^{(|A|+|B|)\times(|A|+|B|)}$ is a diagonal matrix where $\mathbf{D}_{ii}$ represents the count of nonzero elements in the $i$-th row of $\mathbf{W}$. Specifically, for single-domain graphs A and B and the cross-domain graph, we separately define their transition matrices as $\mathbf{\widehat{W}}_{A}$, $\mathbf{\widehat{W}}_{B}$, $\mathbf{\widehat{W}}_{C}$.
Propagation and Aggregation
For each graph, we adopt a widely used LightGCN-based propagation approach, which abandons feature transformation and non-linear activation, instead opting for simple weighted sum aggregation operators, with the $k$-th layer’s propagation defined as
$\mathbf{e}_{i}^{(k+1)}=\sum_{j\in\mathcal{N}_{i}}\frac{\textbf{T}_{ij}}{\sqrt{\left|\mathcal{N}_{i}\right|}\sqrt{\left|\mathcal{N}_{j}\right|}}\mathbf{e}_{i}^{(k)},$ (4)
where $\mathbf{e}^{(k)}_{i}$ is in $\mathbf{E}^{(k)}=[\mathbf{e}^{(k)}_{1};\mathbf{e}^{(k)}_{2};\cdots]$, and $\mathcal{N}_{i}$ and $\mathcal{N}_{j}$ respectively represent the number of neighbors for nodes $i$ and $j$ in the graph. Then we can formulate the matrix equivalent form of the propagation for the three graphs as
$\left\{\begin{array}[]{l}\mathbf{E}^{(k+1)}_{A}=\mathbf{\widehat{W}}_{A}\mathbf{E}^{(k)}_{A},\\ \mathbf{E}^{(k+1)}_{B}=\mathbf{\widehat{W}}_{B}\mathbf{E}^{(k)}_{B},\\ \mathbf{E}^{(k+1)}_{C}=\mathbf{\widehat{W}}_{C}\mathbf{E}^{(k)}_{C}.\\ \end{array}\right.$ (5)
Finally, after passing through $K$ layers, the final embeddings of the three graphs are obtained by:
$\left\{\begin{array}[]{l}\mathbf{E}_{A}=\frac{1}{K}\sum_{k=0}^{K}\mathbf{E}^{(k)}_{A},\\ \mathbf{E}_{B}=\frac{1}{K}\sum_{k=0}^{K}\mathbf{E}^{(k)}_{B},\\ \mathbf{E}_{C}=\frac{1}{K}\sum_{k=0}^{K}\mathbf{E}^{(k)}_{C}.\\ \end{array}\right.$ (6)
Cross-Transition Multi-Head Self-Attention
In this section, we introduce the Cross-Transition Multi-Head Self-Attention module in $\text{Transition}^{2}$ that captures cross-domain transition and feedback transition information by equipping the multi-head self-attention module with a cross-mask mechanism.
Firstly, to model domain transitions and feedback transitions separately in the cross-domain user history, given the user’s ordered interaction history $\{\text{item}_{1},\text{item}_{2},\ldots\}$ with corresponding domain identifiers $\{d_{1},d_{2},\ldots\}$ and feedbacks $\{f_{1},f_{2},\ldots\}$, we first obtain the cross-domain embedded representation $\mathbf{E}_{C}$ through the transition-aware graph encoder for this sequence. Next, to capture sequential relationships, we further introduce positional embedding $\mathbf{P}_{C}$. For the self-attention mechanism, we define $\widehat{\mathbf{E}}_{C}=\mathbf{E}_{C}+\mathbf{P}_{C}$ as the input. However, simply using all embeddings as input without any restrictions would make it difficult for the self-attention mechanism to distinguish between feedback and domain information. Therefore, we introduce two masks $\mathbf{M}^{f}$ and $\mathbf{M}^{d}$ to better capture feedback transitions and domain transitions, with each entry defined as:
$\mathbf{M}^{f}_{ij}=\begin{cases}0&\text{ if }f_{i}=f_{j},\\ 1&\text{ if }f_{i}\neq f_{j}.\end{cases}$ (7)
$\mathbf{M}^{d}_{ij}=\begin{cases}0&\text{ if }d_{i}=d_{j},\\ 1&\text{ if }d_{i}\neq d_{j}.\end{cases}$ (8)
However, simply introducing these two masking mechanisms is not sufficient to strongly capture cross-transition information. Therefore, we perform a cross-processing of these two masks to obtain cross-masks that can capture four types of cross-transition information:
$\left\{\begin{array}[]{l}\mathbf{M}_{1}=\mathbf{M}^{f}\&\mathbf{M}^{d},\\ \mathbf{M}_{2}=\mathbf{M}^{f}\&\widetilde{\mathbf{M}}^{d},\\ \mathbf{M}_{3}=\widetilde{\mathbf{M}}^{f}\&\mathbf{M}^{d},\\ \mathbf{M}_{4}=\widetilde{\mathbf{M}}^{f}\&\widetilde{\mathbf{M}}^{d}.\end{array}\right.$ (9)
where & represents bitwise AND, and $\widetilde{\mathbf{M}}$ represents the negation of $\mathbf{M}$. To account for all four types of cross-transitions, we equip each head of the Multi-Head Self-attention with a different mask. Specifically, we set the overall mask as $\mathbf{M}_{C}=[\mathbf{M}_{1};\mathbf{M}_{2};\mathbf{M}_{3};\mathbf{M}_{4};\cdots]$. The Cross-Transition MHSA then takes $\widehat{\mathbf{E}}_{C}$ and $\mathbf{M}$ as inputs, with $\mathbf{Q}=\mathbf{K}=\mathbf{V}=\widehat{\mathbf{E}}_{C}$, and a Feed-Forward layer following, and output the final embedding $\mathbf{E}^{\prime}_{C}$:
$\mathbf{E}^{\prime}_{C}=\mathrm{FFN}\left(\mathrm{MHSA}\left(\widehat{\mathbf{E}}_{C},\widehat{\mathbf{E}}_{C},\widehat{\mathbf{E}}_{C},\mathbf{M}_{C}\right)\right).$ (10)
Specifically, the detailed computation of Multi-Head Self-Attention is as follows:
$\mathrm{MHSA}\left(\widehat{\mathbf{E}}_{C},\widehat{\mathbf{E}}_{C},\widehat{\mathbf{E}}_{C},\mathbf{M}_{C}\right)=\text{Concate}\left(\text{head}_{1},\text{head}_{2},\ldots\right)W.$ (11)
where for each headi=1,2,…,
$\text{head}_{i}$ $=\mathrm{Attention}\left(\widehat{\mathbf{E}}_{C},\widehat{\mathbf{E}}_{C},\widehat{\mathbf{E}}_{C},\mathbf{M}_{i\%4}\right)$ (12)
$=\mathrm{Softmax}\left(\widehat{\mathbf{E}}_{C}\widehat{\mathbf{E}}_{C}^{\top}/\sqrt{d/h}\odot\mathbf{M}_{i\%4}\right)\widehat{\mathbf{E}}_{C},$
where $\odot$ denotes the Hadamard product.
However, for single-domain sequences, domain transitions do not occur, so we only consider feedback transitions. Specifically, for the single-domain embedded representations $\mathbf{E}_{A}$ and $\mathbf{E}_{B}$ after the Transition-Aware Graph encoder and the positional embedding $\mathbf{P}_{A}$ and $\mathbf{P}_{B}$, we define $\widehat{\mathbf{E}}_{A}=\mathbf{E}_{A}+\mathbf{P}_{A}$, $\widehat{\mathbf{E}}_{B}=\mathbf{E}_{B}+\mathbf{P}_{B}$, $\mathbf{M}_{A}=\mathbf{M}_{B}=[\mathbf{M}^{f};\widetilde{\mathbf{M}}^{f};\mathbf{M}^{f};\cdots]$. Then, the final embedded representations are obtained through a similar Multi-Head Self-Attention mechanism:
$\mathbf{E}^{\prime}_{A}=\mathrm{FFN}\left(\mathrm{MHSA}\left(\widehat{\mathbf{E}}_{A},\widehat{\mathbf{E}}_{A},\widehat{\mathbf{E}}_{A},\mathbf{M}_{A}\right)\right),$ (13)
$\mathbf{E}^{\prime}_{B}=\mathrm{FFN}\left(\mathrm{MHSA}\left(\widehat{\mathbf{E}}_{B},\widehat{\mathbf{E}}_{B},\widehat{\mathbf{E}}_{B},\mathbf{M}_{B}\right)\right),$ (14)
Model Training
For the final model training loss, we designed losses $\mathcal{L}_{\text{single }}^{A}$ and $\mathcal{L}_{\text{single }}^{B}$ for single-domain recommendation tasks and losses $\mathcal{L}_{\text{cross }}^{A}$ and $\mathcal{L}_{\text{cross }}^{B}$ for cross-domain recommendation tasks. These are then combined to obtain the total loss:
$\mathcal{L}_{\text{total }}=\underbrace{\mathcal{L}_{\text{single }}^{A}+\mathcal{L}_{\text{single }}^{B}}_{\text{Single-Domain Loss }}+\underbrace{\mathcal{L}_{\text{cross }}^{A}+\mathcal{L}_{\text{cross }}^{B}}_{\text{Cross-Domain Loss }}.$ (15)
where the losses are all defined as the cross-entropy loss between the predictions and the ground truth:
$\left\{\begin{array}[]{l}\mathcal{L}_{\text{single }}^{A}=-\operatorname{log}\mathrm{Softmax}\left(\mathrm{MLP}_{A}(\mathbf{E}^{\prime}_{C}+\mathbf{E}^{\prime}_{A})\right)_{a_{|S_{A}|+1}},\\ \mathcal{L}_{\text{single }}^{B}=-\operatorname{log}\mathrm{Softmax}\left(\mathrm{MLP}_{B}(\mathbf{E}^{\prime}_{C}+\mathbf{E}^{\prime}_{B})\right)_{b_{|S_{B}|+1}},\\ \mathcal{L}_{\text{cross }}^{A}=-\operatorname{log}\mathrm{Softmax}\left(\mathrm{MLP}_{A}(\mathbf{E}^{\prime}_{C})\right)_{a_{|S_{A}|+1}},\\ \mathcal{L}_{\text{cross }}^{B}=-\operatorname{log}\mathrm{Softmax}\left(\mathrm{MLP}_{B}(\mathbf{E}^{\prime}_{C})\right)_{b_{|S_{B}|+1}}.\end{array}\right.$ (16)
Experiments
Dataset Douban
Domain Book Movie Music
Users $26,877$ $28,718$ $23,822$
Items $187,520$ $57,424$ $185,574$
Records $1,097,148$ $2,828,585$ $1,387,216$
Density $0.0218\%$ $0.1715\%$ $0.0314\%$
Avg. Rating $4.0391$ $3.8101$ $4.1749$
Table 1: Statistics of three domains on Douban.
Methods Book-domain recommendation Movie-domain recommendation
MRR $\mathrm{NDCG}$ $\mathrm{HR}$ MRR NDCG $\mathrm{HR}$
@$10$$5$$10$$1$$5$$10$$10$$5$$10$$1$$5$$10$
GRU4Rec $1.31$ $1.28$ $1.39$ $0.93$ $1.59$ $1.97$ $1.62$ $1.58$ $1.80$ $0.99$ $2.13$ $2.83$
SASRec $1.35$ $1.32$ $1.43$ $0.98$ $1.62$ $1.97$ $1.89$ $1.87$ $2.09$ $1.24$ $2.45$ $3.13$
SRGNN $1.40$ $1.38$ $1.48$ $1.00$ $1.70$ $2.00$ $1.69$ $1.63$ $1.84$ $1.11$ $2.12$ $2.78$
CoNet $1.35$ $1.35$ $1.48$ $0.96$ $1.72$ $2.11$ $1.62$ $1.55$ $1.79$ $1.01$ $2.07$ $2.83$
$\pi$-Net $1.42$ $1.44$ $1.61$ $0.91$ $1.93$ $2.48$ $1.62$ $1.58$ $1.85$ $0.90$ $2.22$ $3.08$
C2DSR $1.52$ $1.49$ $1.64$ $1.06$ $1.89$ $2.45$ $1.93$ $1.92$ $2.15$ $1.25$ $2.54$ $3.26$
TriCDR $\underline{1.55}$ $\underline{1.54}$ $\underline{1.68}$ $\underline{1.10}$ $\underline{1.95}$ $\underline{2.39}$ $\underline{2.00}$ $\underline{1.98}$ $\underline{2.23}$ $\underline{1.30}$ $\underline{2.62}$ $\underline{3.37}$
$\text{Transition}^{2}$ 1.63 1.62 1.77 1.15 2.04 2.51 2.32 2.33 2.59 1.53 3.05 3.80
Table 2: Experimental results (%) of the Book-Movie domains. The best result is bolded and the runner-up is underlined. Improvements over the second-best methods are significant (t-test, p-value $<$ 0.05).
Methods Book-domain recommendation Music-domain recommendation
MRR $\mathrm{NDCG}$ $\mathrm{HR}$ MRR NDCG $\mathrm{HR}$
@$10$$5$$10$$1$$5$$10$$10$$5$$10$$1$$5$$10$
GRU4Rec $1.24$ $1.19$ $1.31$ $0.90$ $1.47$ $1.84$ $1.18$ $1.15$ $1.29$ $0.77$ $1.51$ $1.93$
SASRec $1.22$ $1.20$ $1.30$ $0.88$ $1.49$ $1.82$ $1.25$ $1.23$ $1.36$ $0.84$ $1.59$ $1.98$
SRGNN $1.24$ $1.22$ $1.34$ $0.91$ $1.52$ $1.86$ $1.28$ $1.25$ $1.39$ $0.86$ $1.61$ $2.05$
CoNet $1.30$ $1.31$ $1.44$ $0.90$ $1.69$ $2.11$ $1.23$ $1.21$ $1.39$ $0.77$ $1.64$ $2.18$
$\pi$-Net $1.28$ $1.26$ $\underline{1.45}$ $0.75$ $\underline{1.75}$ 2.35 $1.20$ $1.17$ $1.38$ $0.64$ $1.67$ $\underline{2.34}$
C2DSR $1.30$ $1.26$ $1.38$ $0.94$ $1.56$ $1.92$ $1.30$ $\underline{1.30}$ $\underline{1.43}$ $\underline{0.88}$ $\underline{1.69}$ $2.10$
TriCDR $\underline{1.34}$ $\underline{1.31}$ $1.43$ $\underline{0.95}$ $1.64$ $2.03$ $\underline{1.30}$ $1.28$ $1.42$ $0.86$ $1.67$ $2.10$
$\text{Transition}^{2}$ 1.40 1.39 1.53 0.96 1.80 $\underline{2.22}$ 1.41 1.40 1.57 0.90 1.87 2.38
Table 3: Experimental results (%) of the Book-Music domains. The best result is bolded and the runner-up is underlined. Improvements over the second-best methods are significant (t-test, p-value $<$ 0.05).
We conduct extensive experiments and detailed studies to evaluate the performance of $\text{Transition}^{2}$.
Expermental Setting
Datasets
To verify the effectiveness of the proposed $\text{Transition}^{2}$, we conduct experiments on a public dataset from Douban [^2] https://www.douban.com/. We select two CDSR scenarios for experiments: “Book-Movie” and “Book-Music”. To ensure sequential constraints, we retain cross-domain interaction sequences that include at least 3 items from each domain within one year. Ratings greater than 3 are considered positive feedback, while ratings less than or equal to 3 are considered negative feedback. We preprocess the data into a sequential recommendation format and split it based on timestamp records: the first 80% serves as the training set, the next 10% as the validation set, and the final 10% as the test set. Statistics are provided in Table 1.
Baselines
In this section, We compare $\text{Transition}^{2}$ with three representative sequential recommendation (SR) baselines and five cross-domain SR baselines:
Sequential recommendation baselines:
  • GRU4Rec (Hidasi et al. 2015) uses GRUs to capture sequential patterns in session-based data for personalized recommendations.
  • SASRec (Kang and McAuley 2018) uses self-attention to capture long-range dependencies in user behavior for effective recommendations.
  • SRGNN (Wu et al. 2019) applies Graph Neural Networks (GNN) to session-based recommendation, representing user behaviors as graphs to capture complex relationships and improve accuracy.
Cross-domain sequential recommendation baselines:
  • CoNet (Hu, Zhang, and Yang 2018) models interactions in two domains using base networks and transitions information via a cross-network.
  • $\pi$-Net (Ma et al. 2019) introducing a novel gating recurrent module to model and transfer knowledge across different domains.
  • C2DSR (Cao et al. 2022) uses a graph neural network to leverage inter-domain co-occurrences and employs a contrastive infomax objective to transfer cross-domain preferences by maximizing mutual information.
  • TriCDR (Ma et al. 2024) uses triple cross-domain attention and contrastive learning to model comprehensive cross-domain correlations.
Evaluation Metrics
To ensure unbiased evaluation, we employ the leave-one-out method, consistent with methodologies used in previous studies (Ye, Li, and Yao 2023; Kang and McAuley 2018). Following Rendle’s approach (Krichene and Rendle 2020), each validation/test case is assessed with $1,000$ scores, comprising $999$ negative items and $1$ positive item. The Top-K recommendation performance across these $1,000$ ranking lists is evaluated using metrics such as MRR@$10$ (Mean Reciprocal Rank) (Voorhees et al. 1999), NDCG@$5$, $10$ (Normalized Discounted Cumulative Gain) (Järvelin and Kekäläinen 2002), and HR@$1$, $5$, $10$ (Hit Ratio) (Koren, Bell, and Volinsky 2009).
Implementation Details
Our algorithm is implemented in PyTorch (Paszke et al. 2019) . The embedding size ($D$) and mini-batch size ($N$) are fixed at $256$, with training epochs set to $100$ and dropout fixed at $0.2$. We employ Adam (Kingma and Ba 2014) as the optimizer for parameter updates. The $L_{2}$ regularizer coefficient is chosen from {$-41e\text{-}$ 4, $-55e\text{-}$ 5, $-51e\text{-}$ 5}, and the learning rate $lr$ is selected from {$-31e\text{-}$ 3, $-45e\text{-}$ 4, $-41e\text{-}$ 4}. In C2DSR, we vary the depth of the GNN $L$ from {$1$, $2$, $3$}, and adjust the harmonic factor $\lambda$ from $0.1$ to $0.9$ in increments of $0.1$. For SASRec-based models, we incorporate two single-head attention blocks and learnable position embeddings. The channel number is set to $5$ for $\pi$-net, consistent with the original paper (Ma et al. 2019). In our approach, $N_{H}$ for Head Number is chosen from {$4$, $8$, $12$, $16$}. Best-performing models are determined based on the highest Mean Reciprocal Rank (MRR) performance on the validation set, and their results are reported on the test set. The experimental setup includes Ubuntu 18.04 with an Intel Xeon Gold 5218 CPU (64 cores, 128 threads), 754GB RAM, and four NVIDIA GeForce RTX 3090 GPUs, each with 24GB of memory. The GPU driver version is 550.78, and the CUDA version is 12.4.
Methods Book-domain recommendation Movie-domain recommendation
MRR $\mathrm{NDCG}$ $\mathrm{HR}$ MRR NDCG $\mathrm{HR}$
@$10$$5$$10$$1$$5$$10$$10$$5$$10$$1$$5$$10$
w/o TAG $1.58$ $1.57$ $1.73$ $1.08$ $2.03$ $2.50$ $2.27$ $2.28$ $2.54$ $1.50$ $3.01$ $3.80$
w/o CTM $1.56$ $1.56$ $1.69$ $1.13$ $1.95$ $2.36$ $2.17$ $2.18$ $2.42$ $1.43$ $2.87$ $2.62$
only $\mathbf{M}_{1}$ $1.59$ $1.59$ $1.72$ $1.14$ $2.00$ $2.41$ $2.36$ $2.37$ $2.63$ $1.56$ $3.11$ $3.92$
only $\mathbf{M}_{2}$ $1.56$ $1.56$ $1.69$ $1.12$ $1.95$ $2.36$ $2.16$ $2.16$ $2.41$ $1.41$ $2.85$ $3.62$
only $\mathbf{M}_{3}$ $1.56$ $1.56$ $1.69$ $1.13$ $1.95$ $2.35$ $2.18$ $2.18$ $2.43$ $1.42$ $2.88$ $3.64$
only $\mathbf{M}_{4}$ $1.55$ $1.54$ $1.68$ $1.13$ $1.93$ $2.34$ $2.19$ $2.19$ $2.44$ $1.44$ $2.88$ $3.64$
only $\mathbf{M}_{f}$ $1.57$ $1.57$ $1.71$ $1.14$ $1.97$ $2.39$ $2.31$ $2.32$ $2.58$ $1.51$ $3.06$ $3.78$
only $\widetilde{\mathbf{M}}_{f}$ $1.54$ $1.54$ $1.65$ $1.13$ $1.90$ $2.26$ $2.29$ $2.31$ $2.54$ $1.54$ $3.01$ $3.75$
$\text{Transition}^{2}$ $1.63$ $1.62$ $1.77$ $1.15$ $2.04$ $2.51$ $2.32$ $2.33$ $2.59$ $1.53$ $3.05$ $3.80$
Table 4: Ablation studies (%) on the Book-Movie Domains.
Experimental Results and Analysis
Overall Performance
Table 2 and Table 3 demonstrate the performance of the methods in the “book-movie” and “book-music” CDSR scenarios.
(1) For SR baselines, GRU4Rec, SASRec and SRGNN perform well, with SRGNN performing the best. This validates that modeling interaction sequences using graph neural networks can provide valuable insights for making accurate recommendations.
(2) For CDSR baselines, most baselines perform better than SR baselines because considering knowledge transfer between domains helps improve cross-domain recommendations. CoNet’s performance is slightly lower than SR baselines, possibly due to the weaker capability of its model architecture.
(3) Our $\text{Transition}^{2}$ significantly outperforms all baselines across all metrics, demonstrating the superiority of our model in the CDSR task. These experimental results prove that considering domain transition and feedback transition in modeling single-domain and cross-domain recommendations is highly effective.
Ablation Studies
Figure 3: Sensitivity Analysis of $L$.
Figure 4: Sensitivity Analysis of $N_{H}$.
The results of the ablation study, presented in Table 4 for the ”Book-Movie” scenario, reveal several key insights into the contributions of various components within $\text{Transition}^{2}$:
  • Impact of Transition-Aware Graph Encoder and Cross-Transition MHSA: The removal of the Transition-Aware Graph Encoder (w/o TAG) and the replacement of Cross-Transition Multi-Head Self-Attention with regular MHSA (w/o CTM) result in a notable decline in recommendation performance. This clearly demonstrates the critical role these components play in capturing the transitions between different domains. Their combined absence highlights their synergistic importance in maintaining high performance.
  • Effect of Retaining Individual Transition Masks in Cross-Domain Sequence: When the model is configured to retain only individual transition masks in the MHSA of the cross-domain sequence (only $\mathbf{M}_{1}$, only $\mathbf{M}_{2}$, only $\mathbf{M}_{3}$, only $\mathbf{M}_{4}$), the general trend is a degradation in performance. This indicates that each transition mask contributes distinct information, and isolating them misses important cross-domain transitions. However, an interesting exception is observed with only $\mathbf{M}_{1}$ in the Movie domain, where a slight improvement is noted. This could imply that in certain cases, specific transitions might hold more relevance, but the overall performance gain from retaining all masks indicates that the comprehensive transition information is essential for robust recommendations.
  • Performance with Single-Domain Sequence Masks: In the MHSA of the single-domain sequence, retaining only feedback transition information (only $\mathbf{M}_{f}$) or only non-feedback transition information (only $\widetilde{\mathbf{M}}_{f}$) leads to a reduction in performance. This finding underscores the importance of the mask design in $\text{Transition}^{2}$, which balances both transition and non-transition information. The decline in performance when these elements are isolated suggests that both types of information are integral to accurately modeling user preferences within a single domain, further validating the effectiveness of the dual-mask approach in $\text{Transition}^{2}$.
Hyper-parameters
This section investigates the parameter sensitivity of the depth $L$ of the graph encoder and the number of self-attention heads $N_{H}$. For fair comparison, when studying $L$, we fix $N_{H}=8$; when studying $N_{H}$, we fix $L=1$. For the hyperparameter $L$, we conducted analysis experiments with values in {0, 1, 2, 3}, where $L=0$ means ignoring the Transition-Aware Graph Encoder. Figure 3 shows the impact on recommendation performance in terms of MRR@10 and NDCG@10 metrics for the “Book” and “Movie” domains, where it is observed that our model achieves the best performance when $L=1$ on Movie domain. For the hyperparameter $N_{H}$, we selected values from {4, 8, 12, 16}, and Figure 4 show the impact on recommendation performance in terms of MRR@10 and NDCG@10 metrics for the ”Book” and ”Movie” domains, where $\text{Transition}^{2}$ achieves the best recommendation performance on Book domain when $N_{H}=8$.
Conclusion
In this paper, we proposed $\text{Transition}^{2}$, a novel approach to model transitions across both domains and types of user feedback in cross-domain sequential recommendations. By using a transition-aware graph encoder and cross-transition multi-head self-attention, our method effectively captures and integrates the transition information from user history. Experimental results on two public datasets demonstrate the effectiveness of $\text{Transition}^{2}$ in improving prediction accuracy across different domains.